OpenStudy (anonymous):

Will fan and give medal to whoever helps! two integers, x and y have a product of 20. what is the least possible sum of x and y?

OpenStudy (anonymous):

5 and4

OpenStudy (anonymous):

but least addition is 5*4=20 and 5+4=9 that is least than 10+2=12

Directrix (directrix):

@Distroyaology It is not stated that the two numbers with a product of 20 must be positive. (-20)* (-1) = 20 -20 + (-1) = -21 which appears to be the LEAST possible sum. -->